# Assortment Optimization for Patient-Provider Matching

### v1

@@ 3.5 Worked Example with Policies
    Gradient Descent optimizes for $g(f(\mathbf{X}))\cdot\theta=f(\mathbf{X})\cdot\theta=\sum_{i=1}^{N}X_{i,1}\theta_{i,1}h(\lVert\mathbf{X}_{*,j}\rVert_{1})$, which is optimizes at $\mathbf{X}=[1,1,0]$ and is the optimal assortment for match quality.
The greedy policy fails because it offers the provider to patients with a low match quality, while the pairwise policy fails because it offers providers to too few patients. The group-based and gradient descent policies rectify this situation by adapting assortment size based on the match quality and correctly offering the provider to patients 1 and 2.
4 Extending Policies to Real-World Considerations
We extend our model to incorporate three real-world matching phenomena: 1) batch-offering assortments, 2) patient cognitive load, and 3) fairness considerations.
4.1 Batch Offering Patients
Thus far, we have assumed that all menus are constructed up front, in a single “batch.” Healthcare administrators can sacrifice logistical ease to improve match quality by making offers to groups of patients in batches rather than all at once. By offering assortments in batches, administrators increase control over the response order but increase the total response time for all patients to respond. By assumption, the response time for each patient is exponentially distributed (see Section 2.1), and the expected maximum of $N$ exponential variables is proportional to $\log(N)$ (Eisenberg, 2008). By offering $L$ batches, our total response time increases from $\log(N)$ to $L\log(N/L)$.
We aim to construct batches, $b_{1},b_{2},\ldots,b_{L}\subseteq[N]$, so that the resulting response order maximizes match quality. Formally, let $S(b_{1},b_{2},\ldots,b_{L})$ be a distribution over response orders so that any patient $i\in b_{k}$ is before any patient $i^{\prime}\in b_{k^{\prime}}$ with $k<k^{\prime}$. Next, let $\sigma^{*}$ be the optimal ordering for some fixed policy $\pi$: $\sigma^{*}=\operatorname*{arg\,max}_{\sigma}\frac{1}{N}\sum_{t=1}^{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi(\theta,\sigma)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})$. We aim to set $b_{1},b_{2},\ldots,b_{L}$ so that $\sigma\sim S(b_{1},b_{2},\ldots,b_{L})$ and $\sigma^{*}$ are similar.
To do so, we first characterize $\sigma^{*}$ for a policy $\pi^{A}(\theta,\sigma)$, which augments the pairwise policy using knowledge of $\sigma$, $\pi^{A}(\theta,\sigma)_{i,j}\geq\pi^{P}(\theta)_{i,j}\forall i,j$. Recall that $v_{i}=j$ if $\pi^{P}(\theta)_{i,j}=1$, and let $v^{-1}_{j}=i$ if $v_{i}=j$. Then, $\pi^{A}(\theta)_{i,j}=1$ if $\theta_{i,j}\geq\theta_{i,v_{i}}$ and $\sigma^{-1}_{i}\geq\sigma^{-1}_{v^{-1}_{j}}$, where $\sigma^{-1}_{i}=t$ if $\sigma_{t}=i$. Essentially, $\pi^{A}(\theta)$ augments the pairwise policy $\pi^{P}(\theta)$ to add pairs $i,v_{i^{\prime}}$ for patients $i^{\prime}$ preceding $i$ in $\sigma$. We focus on such a policy because it represents a natural extension of the pairwise policy to incorporate the order $\sigma$. We characterize $\sigma^{*}$ for the policy $\pi^{A}(\theta)$ then use this to construct the batches $b_{k}$:
Lemma 4.1.
Let $G=(V,E)$ be a graph with $N$ nodes such that node $i$ is connected to $i^{\prime}$ if $\theta_{i,v_{i}}\leq\theta_{i,v_{i^{\prime}}}$ or $v_{i}=-1$ and $v_{i^{\prime}}\neq-1$. Let $\sigma^{*}$ be the optimal ordering for the policy $\pi^{A}$. Then, traversing the nodes defined by $\sigma^{*}$ is a reverse topological ordering on $G$.
We demonstrate this by considering a pair of patients $i$ and $i^{\prime}$, so that $i$ precedes $i^{\prime}$ in any reverse topological order, but $i$ precedes $i^{\prime}$ in $\sigma^{*}$. We then use case work to show that we can either a) show that there is a cycle in the graph $G$, leading to a contradiction, b) recurse on a smaller pair of patients, $w,i$ or $i^{\prime},w$, or c) show that swapping $i$ and $i^{\prime}$ in $\sigma$ can only increase the match quality.
We next construct $b_{1},b_{2},\ldots,b_{L}$, so $\sigma\sim S(b_{1},b_{2},\ldots,b_{L})$ is close to the optimal ordering $\sigma^{*}$. The idea is to partition $\sigma^{*}$ into $L$ batches; then we know that all patients in $b_{1}$ precede $b_{2}$. To do this, we first demonstrate that if no pairs of patients within a batch share a common descendant, then such a batching preserves the match quality of $\sigma^{*}$: {theorem}[] Let $\sigma^{*}$ be an optimal ordering. Let $G=(V,E)$ be a graph with $N$ nodes such that node $i$ is connected to $i^{\prime}$ if $\theta_{i,v_{i}}\leq\theta_{i,v_{i^{\prime}}}$ or $v_{i}=-1$ and $v_{i^{\prime}}\neq-1$. Suppose that there exists a partition of $\sigma^{*}$ into $K$ batches, $b_{1},b_{2},\ldots,b_{L}$, such that for any partition $b_{k}$, no $i,i^{\prime}\in b_{k}$ have a common descendant in $G=(V,E)$. Then $\pi^{A}$ achieves the same match quality under partition $g$ as under the optimal ordering:
$\mathbb{E}_{\sigma\sim S(b_{1},b_{2},\ldots,b_{L})}[\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma_{t}}\left(\pi^{A}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})]=\frac{1}{N}\sum_{t=1}^{N}(f_{\sigma^{*}_{t}}\left(\pi^{A}(\theta)_{\sigma_{t}}\odot\mathbf{y}^{(t)}\right)\cdot\theta_{\sigma_{t}})$ (19)
This holds because the selection made by patient $i$ depends only on the descendants of patient $i$ on the graph G. If $b_{1},b_{2},\ldots,b_{K}$ ensures that any pair of descendants are in different batches, then the order of descendants is maintained, so patients select the same providers in $\sigma$ and $\sigma^{*}$.
In practice, we compute the batches $b_{k}$ through the following:
  1. 1.
    Computing the graph $G=(V,E)$ with $i$ connected to $i^{\prime}$ if $\theta_{i,v_{i^{\prime}}}>\theta_{i,v_{i}}$.
  2. 2.
    Computing a reverse topological ordering $\sigma^{*}$.
  3. 3.
    Selecting batches with the minimum number of edges within a batch (via dynamic programming).
We note that this is an approximation of optimal batches for two reasons: 1) while $\sigma^{*}$ is a reverse topological ordering, there could be multiple reverse topological orderings, and 2) minimizing the number of edges within a batch does not guarantee that the order of descendants is maintained.
4.2 Maximum Assortment Sizes
To incorporate patient cognitive load (Chernev et al., 2015), we incorporate restrictions on the maximum assortment size into our model. Large assortments might be undesirable or practically infeasible because patients only consider a subset of the options they are presented due to logistical or time constraints. Formally, we suppose that patients only consider a random $d$-subset $\mathbf{z}^{(d)}$ of the options they are presented at random, $z^{(d)}_{j}\leq X_{i,j}y^{(t)}_{j}$ with $\lVert\mathbf{z}^{(d)}\rVert_{1}\leq d$. Patients make decisions according to $f_{i}(\mathbf{z}^{(d)})$ rather than $f_{i}(\pi(\theta)\odot\mathbf{y}^{(t)})$. When $d$ is small, policies should offer small tailored assortments, as large assortments lead to patients ignoring many options.
4.3 Fair Matches
Policies should aim to achieve fair matches to encourage participation and maintain equity. We construct examples of fairness metrics through the prior literature on fair matches and list these below (Xinying Chen and Hooker, 2023; García-Soriano and Bonchi, 2020).
  • Variance in Match Quality - Another approach minimizes variance in match quality to ensure that the discrepancy in match quality between patients is small.
Each notion of fairness has different implications for the optimal policy. For example, optimizing for the minimum match quality might lead policies to only offer assortments to high match quality patient-provider pairs, whereas minimizing the variance in match quality might yield many low quality matches. We explore tradeoffs in match quality and fairness in Section 5.4.
5 Empirical Analysis
We analyze policies in synthetic and real-world settings to characterize how assortment policy impacts match quality and rate. First, we demonstrate that gradient descent is the best policy when patients outnumber providers. We then show that group-based and greedy policies perform well for large and small values of match probability $p$, respectively. Our results demonstrate that assortment policies should be selected based on problem specifics.
5.1 Experimental Details
We compare the four policies from Section 3 (greedy, pairwise, group-based, and gradient descent) and a random baseline with $\pi(\theta)_{i,j}\sim\text{Ber}(\frac{1}{2})$. For a fixed problem setup, we randomly sample $T$ permutations of patients and compute the match quality and match rate for the corresponding patient order ($\sigma$). We average runs across 15 seeds and $T=100$ trials and plot the standard error across these seeds. Because the scale of match quality and rate might vary between experiments, we report the normalized match quality and match rate (norm. MQ and norm. MR for short), which divides each quantity by that of the random policy.
5.2 Synthetic Examples
We construct synthetic examples to gain insight into our policies across different scenarios. For each example, we specify values for $M$, $N$, $f_{i}$, and $\theta$.
Example 4.
Figure 2: We find the optimal policy and compute the approximation ratio for match quality compared to optimal. While all non-random policies achieve approximation ratios $\geq 0.9$ when $M\geq N$ (left), the pairwise policy performs poorly and achieves a low approximation ratio when $N\geq M$ (right).
Few patients and providers - We first analyze policies with a small number of patients $N$ and number of providers $M$. Small $N$ and $M$ allow us to compare to the optimal policy $\pi^{*}$ because we can compute $\pi^{*}$ by enumerating all $2^{NM}$ assortments and selecting the one that maximizes match quality. We run two experiments: the first fixes $N=2$ while varying $M$ from $2$ to $5$, and the second fixes $M=2$ while varying $N$ from $2$ to $5$. For both, we uniformly distribute $\theta$, $\theta_{i,j}\sim U(0,1)$, and let the choice model be uniform choice with $p=1/2$. We compute the match quality approximation ratio to understand how close each policy is to the optimal.
In Figure 2, we demonstrate that when $M\geq N$ (left), all non-random policies perform well and achieve within 3.5% of optimal, while for $N\geq M$ (right), the pairwise policy performs poorly. When $M\geq N$, providers are abundant so each patient can match with their most preferred provider. However, when $N\geq M$, the pairwise policy performs up to 28% worse than optimal. Providers are scarce in this scenario, so policies should offer the same provider to multiple patients. The pairwise policy only offers each provider one patient, which is suboptimal and leads to poor match quality.
Example 5.
Figure 3: When there are more patients than providers, the gradient descent policy performs best, both for uniformly and normally distributed $\theta$.
Varied patient/provider ratio - In real-world situations, patients outnumbers providers (HHS, 2013), so we analyze how the patient/provider ratio impacts policy performance. We fix the number of providers $M=25$, while varying the number of patients $N$ from $25$ up to $200$. We compare policies under two match quality distributions: the first uniformly distributes $\theta\sim U(0,1)$, while the second normally distributes $\theta_{i,j}\sim\mathcal{N}(\mu_{j},0.1)$. The former corresponds to heterogeneous preferences where patient preferences are non-correlated, while the latter corresponds to homogenous preferences, where patient preferences are influenced by some signal of provider quality $\mu_{j}$. For both experiments, we fix $p=0.5$ for the uniform choice model.
In Figure 3, we demonstrate that the gradient descent policy performs best when patients outnumber providers. When $N/M=8$, the gradient descent policy performs 5% better than alternatives for uniform $\theta$ and 9% for normal $\theta$. The gradient descent policy performs best for large $N/M$ because $g(f(\mathbf{X}))$ better approximates match quality when $N>M$; see Section 3.4 for details. Both greedy and group-based policies perform similarly, while the pairwise policy performs worst, mirroring trends from Example 4. In Appendix A, we similarly compare policies according to the match rate and find that the gradient descent and greedy policies both maximize the match rate.
Example 6.
Figure 4: The greedy policy performs poorly when $p=0.9$ compared to smaller $p$, for both uniform and normal $\theta$, while the pairwise policy performs worse for smaller $p$, especially when $\theta$ is normally distributed. The group-based policy gets around these issues and performs well when $p$ is large and small.
Impact of $p$ and $\theta$ We assess the impact of match quality $\theta$ and match probability $p$ on policy performance with balanced patient and provider counts. We fix $N=M=25$, vary $p\in\{0.1,0.25,0.5,0.75,0.9\}$, and distribute $\theta$ either uniformly or normally.
In Figure 4, we demonstrate that the group-based policy achieves high match quality. As noted in Proposition 3.1, the group-based policy can improve upon the pairwise policy when $p$ is small. The pairwise policy performs especially poorly when $\theta$ is normally distributed because patient preferences are homogeneous, as all patients prefer some provider $j^{*}=\operatorname*{arg\,max}_{j}\mu_{j}$. The pairwise policy only places $j^{*}$ in one patient’s assortment, whereas $j^{*}$ should optimally be in multiple assortments due to its preferability. We also note the greedy and group-based policies perform well across $p$ and $\theta$; both policies are always within 6% of the best-performing policy. In Appendix A, we demonstrate that policies achieve similar match rates across $\theta$ and $p$.
(a) Uniform
(b) Normal
Figure 5: We characterize which policy maximizes match quality when varying patient/provider ratio and match probability $p$ when $\theta$ is uniformly (left) and normally (right) distributed. The group-based policy performs best for large $p$, the greedy policy performs best for small $p$, and the gradient descent policy performs best when the patient/provider ratio is large.
To understand which policy performs best across $p$, $\theta$, and $N/M$, we plot the policy which maximizes match quality when varying $p\in\{0.1,0.3,0.5,0.7,0.9\}$ and $N/M\in\{1,2,3,4,5,6,7,8\}$ with $M=10$, and distributing $\theta$ uniformly (Figure 5a) and normally (Figure 5b). The gradient descent policy performs best for large patient/provider ratios, which matches the results from Example 5. The group-based policy performs well for small $p$ because the problem more closely resembles bipartite matching (see Theorem 9), while the greedy policy performs well for small $p$ because it becomes advantageous to offer larger assortments (see Proposition 3.1). We finally note the random policy performs best when $p=0.1$ and $N/M=1$ because policies have high variance in that situation, making it difficult to ascertain which policy performs best.
Example 7.
Varied Choice Model - We evaluate the impact of choice models on policy performance by investigating the MNL and threshold choice models. We fix $N=M=25$, let $\theta$ be distributed uniformly, and vary $\gamma\in\{0,0.1,0.25,0.5\}$ for the MNL choice model and vary $\alpha\in\{0,0.1,0.25,0.5,0.75\}$ with $p=0.5$ for the threshold choice model. .
Figure 6: We compare policy performances when varying the exit option $\gamma$ for the MNL choice model, and the threshold $\alpha$ for the threshold choice model. For the MNL choice model, when $\gamma$ is small, the greedy policy performs poorly, while for large $\gamma$, all non-random policies perform similarly. For the threshold choice model, all policies perform similarly across choices of $\alpha$.
In Figure 6, we show that $\gamma$ determines the performance of the greedy policy for the MNL choice model, while policies perform similarly across $\alpha$ for the threshold choice model. As noted in Section 3.2, the choice of $\gamma$ is inversely related to the setting of $p$, with small $\gamma$ analogous to large $p$. Similar to the trends in Figure 4, we find that the greedy policy performs 22% worse than the best policy when $\gamma=0.1$. For the threshold choice model, $\alpha$ has little impact on the best policy, as greedy, pairwise, and gradient descent perform similarly across $\alpha$. In Appendix A, we explore variants of the uniform choice model where the match probability $p$ is misspecified.
5.3 Real-World Dataset
We work with a partner healthcare organization to assess the performance of our policies in a simulation that mimics situations faced in practice.
5.3.1 Constructing Data from Healthcare Providers
We work with a large healthcare provider in Connecticut to develop a simulation that reflects their system dynamics. We select values of $N$, $M$, $\theta$, and $f_{i}$:
@@ 5.3.1 Constructing Data from Healthcare Providers
    Choice Model $f_{i}$ We select the choice model based on previous work which shows that patients are low-effort decision makers who primarily make decisions based on geographic proximity (Salisbury, 1989; Buzza et al., 2011). We incorporate these two factors and let $f_{i}$ be a threshold model, with $\alpha$ representing the maximum distance patients are willing to travel.
@@
    Match Quality $\theta$ We construct $\theta$ to reflect geographic proximity and the presence of comorbidities. Geographic proximity is the top factor impacting patient match quality (Yen and Mounts, 2013; Buzza et al., 2011), while patients with comorbidities are best served by providers with the corresponding specialized training. Formally, for a patient $i$ and provider $j$, let $d_{i,j}$ be the distance between the patient and provider, and let $\beta_{i,j}$ denote whether patient $i$’s comorbidity and $j$’s specialty match. For example, if patient $i$ has a heart-related comorbidity and provider $j$ has a cardiology speciality, then $\beta_{i,j}=1$. We manually match patient comorbidities to provider specialties. We therefore let match quality $\theta_{i,j}=\alpha+(1-\alpha)(\Delta\beta_{i,j}+(1-\Delta)(\frac{\bar{d}}{d_{i,j}}-1)$. Here, $\Delta$ weights between comorbidities and distance and $\alpha$ sets the threshold match quality for a patient at distance $d_{i,j}=\bar{d}$.
5.3.2 Real-World Simulation Results
@@
Figure 7: (a) We construct a real-world scenario with $N=1225$ and $M=700$. We find that the gradient descent policy performs best because of the large patient/provider ratio. (b) The gradient descent policy achieves a higher match quality for patients with comorbidities because these patients match with specialist providers. (c) Geographically, the gradient descent policy tends to match patients from more populated zip codes, (d) with the highest match rates occurring in cities such as Hartford and New Haven.
@@
In Figure 7a, we show that the gradient descent policy achieves the highest match quality and outperforms alternatives by 13%. This mirrors Figure 5, where gradient descent is the best policy for large patient/provider ratios. Next, we analyze which patients achieve high-quality matches under the gradient descent policy. We find that patients with comorbidities achieve higher match quality ($p<10^{-17}$) because they tend to match with specialist providers (Figure 7b). Additionally, patients from more populated zip codes achieve higher match rates (Figure 7c), potentially due to the lack of providers in rural areas (Hart et al., 2002). Pictorially, this corresponds to high match rates in urban areas such as Hartford and New Haven (Figure 7d).
5.4 Real-World Considerations
We extend our scenario to explore the real-world considerations introduced in Section 4: patient batching, maximum assortment sizes, and fairness considerations.
5.4.1 Batched Patients
(a) Batch Size
(b) Max Assortment Size
(c) Fairness
Figure 8: (a) Offering patients in two batches improves match quality by 7% while increasing response time by 90%. Beyond two batches, the increases in match quality are diminishing. (b) We simulate cognitive overload by assuming patients consider up to $d$ options. Smaller $d$ favors the gradient descent policy because it offers carefully tailored assortments without being too large. (c) While the gradient descent policy maximizes match quality, it achieves worse fairness because it neglects the worst-off patient. In contrast, pairwise and group-based policies achieve higher fairness by only suggesting high-quality matches.
Offering assortments to batches of patients rather than one-shot allows administrators to exercise some control over patient response orders. We apply the batching algorithm from Section 4.1 and analyze the impact of the number of batches, $L$, on the match quality in our real-world scenario. In Figure 8a, we demonstrate that batch offerings allow administrators to trade off logistical burdens for match quality. While offering assortments in two batches can improve match quality by 8% compared to the single-batch setting, the total response time also increases by 74%. Moreover, larger batch sizes lead to diminishing increases in match quality with a near linear increase in the time taken.
5.4.2 Varied Assortment Sizes
Patients might only consider a subset of providers offered due to choice overload (Chernev et al., 2015). To simulate this, we incorporate the maximum assortment size, $d$, to simulate choice overload (see Section 4.2 for details). In Figure 8b, we vary $d\in\{5,10,15,20,25\}$ and find that the gradient descent policy is the best policy across values of $d$, especially for small $d$. This occurs because the gradient descent policy offers smaller assortments compared to the greedy policy, and so restricting assortment size has less of an impact.
5.4.3 Fairness Considerations
To understand how policies trade-off between fairness and match quality, we plot the minimum match quality against the average match quality in Figure 8c (see Section 4.3 for a more thorough discussion of different notions of fairness). In our fairness computation we exclude patients who fail to match, because including them would make the fairness metric always 0 and would not reflect the difference in fairness across policies. We find that the fairest policy, pairwise, also achieves the lowest average match quality. Fairer policies maximize tend to match fewer patients, which keeps the minimum match quality high but lowers the average match quality. If healthcare administrators are focused on fairness as opposed to match quality, then they can select a fairer policy such as the pairwise policy. We compare policies across one definition of fairness for simplicity, but we note that in reality, there are multiple dimensions of fairness that need to be considered (see Section 4.3).

### v2

@@ 4.3 Relaxing Model Assumptions
    Multi-Patient Capacity - We relax the unit capacity assumption for policies by letting policies have a per-provider capacity, $c_{j}=2$. We extend each policy to account for this; the modified policies are detailed in Appendix B. At a high level, each policy has a natural extension to the non-unit capacity scenario; for example, the pairwise policy requires a modification to the constraints of the linear program to account for this.
Figure 6: We explore policy performance when relaxing three assumptions: we experiment with: policies that update online rather than being computed offline, response orders that are proportional to $\theta$ and are not uniformly random, response orders that are controllable by batching patients, and scenarios where providers have non-unit capacities. Across all three relaxations, we have the same ordering of policies regardless of the underlying assumptions. We note that increasing the provider capacity results in homogenization because all policies increase match quality due to increased provider availability.
We evaluate the performance of policies under these relaxed assumptions when $\theta$ is uniformly distributed, and when $f$ is the uniform choice model for $p\in\{0.1,0.25,0.5,0.75,0.9\}$, while fixing $M=25$ and $N=100$. In Figure 6, we find that relaxing assumptions minimally impact policy performance. Under all relaxations, the relative ordering of policies, given $p$, remains the same. Dynamic policies perform better than the pairwise policy, but despite their increased flexibility, are still worse than alternatives such as greedy and gradient descent. Similarly, even policies that can control the order $\sigma$ fail to improve performance compared to the gradient descent policy. Finally, increasing capacity leads to improved performance for the greedy policy, which occurs because provider scarcity is reduced and is equivalent to reducing the ratio of patients to providers (we demonstrate the latter claim through additional experiments in Appendix C). In addition to these experiments, in Appendix A, we demonstrate that our results are robust to misspecification in $\theta$ and $p$.
5 Analysis in a Semi-Synthetic Dataset
We analyze policy performance in a semi-synthetic dataset and show that the gradient descent policy can improve match quality by 13%. We then discuss tradeoffs between various metrics, including match quality, match rate, and fairness.
5.1 Constructing a Semi-Synthetic Simulation
Beyond synthetic data, we work with a large healthcare provider in Connecticut to develop a simulation that reflects their system dynamics, leveraging real-world Medicare data and clinically grounded model parameters. We similarly average policy performance over $T$ rounds with randomized $\sigma$, but select the parameters $N$, $M$, $\theta$, and $f$ to accurately reflect their reality:
@@ 5.1 Constructing a Semi-Synthetic Simulation
    Choice Model $f$ We select the choice model based on previous work which shows that patients are low-effort decision makers who primarily make decisions based on geographic proximity (Salisbury, 1989; Buzza et al., 2011). We incorporate these two factors and let $f$ be a threshold model, with $\alpha$ representing the maximum distance patients are willing to travel.
@@
    Match Quality Matrix $\theta$ We construct $\theta$ to reflect geographic proximity and the presence of comorbidities. Geographic proximity is the top factor impacting patient match quality (Yen and Mounts, 2013; Buzza et al., 2011), while patients with comorbidities are best served by providers with the corresponding specialized training. Formally, for a patient $i$ and provider $j$, let $d_{i,j}$ be the distance between the patient and provider, and let $\beta_{i,j}$ denote whether patient $i$’s comorbidity and $j$’s specialty match. For example, if patient $i$ has a heart-related comorbidity and provider $j$ has a cardiology specialty, then $\beta_{i,j}=1$. The match quality is $\theta_{i,j}=\alpha+(1-\alpha)(\Delta\beta_{i,j}+(1-\Delta)(\frac{\bar{d}}{d_{i,j}}-1)$, where $\Delta$ and $\alpha$ are hyperparameters that modulate the impact of distance and comorbidities. Here, $\Delta$ weighs between comorbidities and distance, and $\alpha$ sets the threshold match quality for a patient at distance $d_{i,j}=\bar{d}$. We provide details on our selection of hyperparameters in Appendix E.
5.2 Dataset Analysis
Figure 7: We compare the distribution of patients (left), providers (middle), and the ratio of providers to patients (right). We find disparities between the distribution of patients and providers, revealing inherent difficulties in designing matching algorithms.
To better understand our semi-synthetic dataset, we analyze the distribution of providers and patients. In Figure 7, we show that the distribution of providers is more clustered in large cities, such as Hartford and New Haven, compared with the distribution of patients, matching results from prior work (Chatterjee, 2022). Moreover, plotting the ratio of providers to patients shows that most zip codes lack any providers while a few near metropolitan areas have a high density of providers (potentially corresponding to healthcare facilities). The distribution of patients and providers reveals inherent tensions for algorithm design; maximizing the match quality might lead to prioritization of “easy-to-match” patients near cities at the expense of more rural patients.
5.3 Match Rate and Match Quality Analysis
@@
Figure 8: (a) We construct a semi-synthetic scenario with $N=1225$ and $M=700$. We find that the gradient descent policy performs best because of the large patient/provider ratio. (b) The gradient descent policy achieves a higher match quality for patients with comorbidities because these patients match with specialist providers. (c) Geographically, the gradient descent policy tends to match patients from more populated zip codes, (d) with the highest match rates occurring in cities such as Hartford and New Haven.
@@
In Figure 8a, we show that on the semi-synthetic dataset, the gradient descent policy achieves the highest match quality and outperforms alternatives by 13%. This mirrors Figure 3, where gradient descent is the best policy for large patient/provider ratios. Next, we analyze which patients achieve high-quality matches under the gradient descent policy. We find that patients with comorbidities achieve higher match quality ($p<10^{-17}$) because they tend to match with specialist providers (Figure 8b). Additionally, patients from more populated zip codes achieve higher match rates (Figure 8c), due to the aforementioned lack of providers in rural areas (Hart et al., 2002). Pictorially, this corresponds to high match rates in urban areas such as Hartford and New Haven (Figure 8d). Our results demonstrate that even if a policy improves performance, it can still result in inequities on who gets matched, and so practitioners should be aware of this when selecting assortment policies.
5.4 Metrics Beyond Match Rate and Match Quality
We next explore policy performance beyond match quality and match rate by introducing metrics such as fairness and regret.
5.4.1 Fairness
Fairness is critical to patient-provider matching to ensure equitable outcomes. To better understand this, we first introduce several definitions of fairness, each inspired by prior work from the fair matching literature (Xinying Chen and Hooker, 2023; García-Soriano and Bonchi, 2020).
  • Variance in Match Quality - Another approach minimizes variance in match quality to ensure that the discrepancy in match quality between matched patients is small.
Each notion of fairness has different implications for the optimal policy. For example, optimizing for the minimum match quality might lead policies to only offer assortments to high match quality patient-provider pairs, whereas minimizing the variance in match quality might yield many low-quality matches.
Figure 9: We compare the performance of policies of both match rate (top) and match quality (bottom) and three definitions of fairness: minimum (left), variance (middle), and range (right) of match quality. We find that a) the three definitions of fairness are aligned, as policies tend to do well or poorly on all three, and b) there are tradeoffs between match rate, match quality, and fairness.
To understand how policies trade-off between match rate, match quality, and fairness, we plot each definition of fairness against both match rate (top) and match quality (bottom) in Figure 9. We find that the three definitions of fairness are well aligned; across all definitions, the pairwise policy performs best, as it maximizes the minimum match quality, and minimizes the range and variance of match qualities. Similarly, the greedy and gradient descent policies perform poorest across all definitions of fairness, though gradient descent performs better than greedy for variance. We additionally find that policies that perform better at either match rate or match quality tend to perform poorer at these fairness metrics. Intuitively, increasing the match quality or match rate requires matching patients with less suitable (lower match quality) providers, which leads to less fair outcomes but improved overall performance. Our overall takeaway is that no one policy performs best across all metrics, so practitioners need to carefully weigh metrics to determine which is best.
5.4.2 Patient Regret
The assortment-based matching process can lead to patient regret if their initial assortment differs from the assortment offered during matching time. While not the primary driver of patient satisfaction, regret can nonetheless impact patient happiness because initially offered providers might no longer be available. To better understand patient regret, we first define it as the difference in the maximum match quality between the original assortment and the offered assortment, where the latter removes providers selected by patients. Formally, we can write regret for patient $i$ as $\max\mathbf{X}_{i}\odot\theta_{i}-\max\mathbf{X}_{i}\odot\theta_{i}\odot\mathbf{y}_{\sigma^{-1}_{i}}$. Here, $\mathbf{y}_{\sigma^{-1}_{i}}$ is the set of available providers when patient $i$ makes a decision.
Figure 10: We compare the regret (left) and assortment size (right) across policies and find that tailoring policies such as through gradient descent or pairwise can reduce regret. Moreover, we find a correlation between assortment size and regret because smaller assortments can allow for less overlap between patients.
To understand how policies perform for regret, in Figure 10, we plot both patient regret and the average assortment size across policies. We find that policies such as pairwise suffer no regret because they offer disjoint menus across patients, while policies such as greedy suffer from the most regret because they do not tailor menus. We also see a general trend between assortment size and regret; as assortment sizes get larger, patient regret tends to increase as well, though the trend is imperfect. Taken together, we find that no policy dominates across all metrics. Improving the match quality might sacrifice fairness or lead to geographical disparities, while policies that perform well for fairness might lead to poor match rates or match qualities in some scenarios.